OpenStudy (kamille):

hey,I need some help solve the equation \[f(x)=f(\frac{ 1 }{ x })\],when\[f(x)=\frac{ x+1 }{ x ^{2} }\] So what I do: \[f(\frac{ 1 }{ x) }=(\frac{ 1 }{ x }+1):\frac{ 1 }{ x ^{2} }=(\frac{ 1+x }{ x }):\frac{ 1 }{ x ^{2} }=x+x^{2}\] then\[\frac{ x+1 }{ x ^{2} }=x+x ^{2}\] and then I got a bit confused, because i dont know how to solve\[x+1+x ^{3}+x ^{4}=0\] Any help?

OpenStudy (kamille):

@amistre64 could you please have a look?

OpenStudy (kamille):

@ParthKohli hey little reddit lover, could you have a look?

OpenStudy (kamille):

hey @Lyrae

OpenStudy (kamille):

dont leave me:(

OpenStudy (kamille):

oo,someone!

OpenStudy (lyrae):

Okay, so there are some complicated ways of solving equations like these, using advenced formulas and stuff. But when you get them at this level there are usually som simple soulution and the simplest way to find one is do some "qualified gusing". If you think of some small integears, are there one which might be a soultion?

OpenStudy (amistre64):

\[f(\frac1x)=\frac{ \frac1x+1 }{ \frac1{x^2} }\] \[=\frac{x^2}{x^2}\frac{ \frac1x+1 }{ \frac1{x^2} }\] \[=\frac{ \frac{x^2}x+1 }{ 1 }\] \[=x+1~:~x\ne0\]

OpenStudy (amistre64):

other than that, im not sure what it is your asking about

OpenStudy (kamille):

well, since the answer is 0, 1 must be eliminated, so 0 can't be true (also because multiplying by 0 is not 'allowed')

OpenStudy (kamille):

so my guess is 1 or -1

OpenStudy (kamille):

(because 1 must be eliminated)

OpenStudy (amistre64):

f(x) = f(1/x) solve for x

OpenStudy (kamille):

yes

OpenStudy (amistre64):

\[x+1=\frac{\frac1x+1}{x^2}\] \[x^2(x+1)=\frac1x+1\] \[x~x^2(x+1)=1+x\]

OpenStudy (amistre64):

well, when x^3 = 1 seems likely

OpenStudy (lyrae):

If you then find a solution you divide the polynomial by that soultion to get one with a lower order. Repeat this until you can solve using quadratic formula.

OpenStudy (kamille):

I dont understand:/

OpenStudy (amistre64):

who dont you understand?

OpenStudy (kamille):

both of you - I didnt understand Lyrae's first message, but not the second one and I dont understand where x^3 dissapears in your equation

OpenStudy (amistre64):

hard to type up all that coding again .... can you retype the process i did and show me were it goes off for you?

OpenStudy (kamille):

alright

OpenStudy (kamille):

I will draw,oke?

OpenStudy (amistre64):

k

OpenStudy (kamille):